OpenStudy (anonymous):

simplify sqrt(x+5)^12

OpenStudy (anonymous):

\[(\sqrt{x+5})^{12}=(x+5)^{\frac{1}{2}\cdot12}=(x+5)^6\]

OpenStudy (anonymous):

(x+5)^6

OpenStudy (anonymous):

(x+5)^6